The distribution in which values of median, mean and mode are equal is considered as: Experimental...

  1. The distribution in which values of median, mean and mode are equal is considered as:

  1. Experimental distribution
  2. Asymmetrical distribution
  3. Symmetrical distribution
  4. Exploratory distribution

  1. Which of the following are true about the Paasche price index?

  1. The index formula uses the weights of the base period
  2. The index formula uses the weights of the current year
  3. The index formula uses the average of the weights of several years
  4. None of the above

  1. The nominal interest is 5 percent, and the annual inflation is 7 percent, then the real interest is
  1. -2%
  2. -20
  3. 12%
  4. 2%
  5. None of the above, we don’t have enough information to answer the question

  1. Which of the following is not a probability-based sample selection method?
    1. Cluster sampling
    2. Stratified sampling
    3. Snowball sampling
    4. Simple Random sampling
  1. According to the central limit theorem, the sampling distribution of the mean can be approximated by the normal distribution:
    1. As the number of the samples gets large enough
    2. As the size of the sample standard deviation decreases
    3. As the sample size (number of observations) gets “large enough”
    4. As the size of the population standard deviation increases

  1. If there is a very strong correlation between two variables then the correlation coefficient must be
    1. Any value larger than 1
    2. Much smaller than 0, if the correlation is negative
    3. Much larger than 0, regardless of whether the correlation is negative or positive
    4. None of these alternatives is correct
  1. In simple linear regression, what does the slope (b1) represent?
    1. The predicted value of Y, when X=0
    2. The estimated change in average Y per unit change in X
    3. The predicted value of Y
    4. The variation around the line of regression
  2. In Linear Regression, what information is contained in the coefficient of determination?
    1. The coefficient of correlation is larger than one.
    2. Whether r has any significance.
    3. We should not partition the total variation.

The proportion of total variation in Y that is explained by X.
Answer #1

1)

The distribution in which values of median, mean and mode are equal is considered as:

  1. Symmetrical distribution

2) The index formula uses the weights of the current year

3)

Real interest rate=[(1+nominal interest rate)/(1+inflation rate)]-1 = (1.05)/(1.07)-1= -1.87% ≈ -2%